OpenStudy (anonymous):

Please help! Find the forth term of the sequence a_n=5-n^2

OpenStudy (anonymous):

We want to evaluate a_4 or where n = 4. All we have to do is plug it into the equation. So... a4=5-4^2 a4=5-16 a4=-11

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

Medal please lol

OpenStudy (anonymous):

I gave you one...

OpenStudy (anonymous):

Alrite alrite jeez lol

OpenStudy (anonymous):

lol